Ax +y = 13
3x -By = -5

Answer by josmiceli(6786)

Ax +y = 13
3x -By = -5
multiply both sides of the 1st equation by B
ABx + By = 13B
3x - By = -5
add the equations
(AB + 3)x = 13B - 5
x = (13B - 5)/(AB + 3)
substitute in the 2nd equation
3(13B - 5)/(AB + 3) - By = -5
By = 3(13B - 5)/(AB + 3) + 5
y = [3(13B - 5)/(AB + 3) + 5]/B